use partial fractions to find the integral

16. x2 - x + 9 J (x2 + 9)2 dx
Answer #1

ANSWER

soli given that By using partial fraction xta cara} cartag arata Cataja zrna + iş Cartag arata (x +90) (art95% + (Aga+ang) Caanci ay =0 .. 998 +92 +9 -0 99, +az -.-- *. from egri 964)+az -9. . q + ₂ =9 =) 92-8-4 = 0 from can ☺ qa, taza-1 910)+932-1 01 D? ch tar auta (x +95 249 29 =o btu -- +9 btl xr-x+9 +9 to [log cart 9) - log crazy11 ) 1. ๆ C * *7) J - น +4 ) ส่ง (7) - (4 +41